OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x − 1) = 5x − (x + 10) explain the process for solving this equation

OpenStudy (anonymous):

remember your order of operations. Do the parenthesis first.

OpenStudy (anonymous):

there, i got cho started.

OpenStudy (anonymous):

Remove the brackets both sides. That will give you 4+(2*x)-(2*1)=5x-(1*x)-(1*10) 4+2x-2=5x-x-10 When simplify 2+2x=4x-10 Now bring all terms containing x to one side. keeping the numbers in the other side. 2+10=4x-2x 12=2x then divide both sides by 2 12/2=(2x)/2 6=x

OpenStudy (rc94):

Something along the lines of: 1) Multiply out brackets 2) Collect like terms on either side 3)Collect like terms on BOTH sides combined and then you'll have the answer

OpenStudy (anonymous):

6=x is the answer?

OpenStudy (rc94):

x=6 is correct. To check you can just put that back into the equation to see if both sides are equal

OpenStudy (anonymous):

whoops sorry.

OpenStudy (anonymous):

its cool skay and thank you RC

OpenStudy (anonymous):

yes Pinkiepie 6 is the answer. to check you can simply plug in 6 for x and if you are correct you will get the same answer on both sides of the equation.